- Table of contents
- Production Status
SAM Expert Task Force
Nate Mayer, Gavin Davies, Adam Aurisano, Andrew Norman, Yuri, Eric Flumerfelt, Matthew Tamsett, Ruth Toner, Kanika Sachdev, Jan Zirnstein, Dominick Rocco, Michael Geith, Chris Backhouse
CD: Marc Mengel, Robert Illingsworth
Last Updated: March 4th 2014
All dates are tentative and optimistic, thus will change organically. They are meant as a guideline for developers only!
February 2014 MC Generation/FarDet Data processing¶
MC generation (first time w/ SAM) and data processing through SAM. Tests ongoing for PID and CAF steps.
SAM Task Force meet Mondays and Thursdays at 11:30am CT to discuss progress and outstanding issues.
Stumbled into a number of issues that have hindered progress. Notably using run numbers starting at 1000000 tripped up on the fcl metadata parser which converted this into a floating point notation (1e6). Solved by checking for int's and then strings in the parser. Issue logged as 5476
Final tests on-going.
Status of tests and metrics can be found here:February 2014 Production
Also metadata module seg faulting at the end of all jobs when trying to close the file but consequently fixed also (backported fixes to tag S14-02-05).
S14-02-05 - FD Cosmics, v2
S14-02-05a - FD genie, v3 (needs to be rerun with a fix to slimming), ND overlay tests
S14-03-06 - PCHitLists, Reconstruction validation
S14-03-20 - FD genie re-run, ND overlays, FD Hadronic cosmic special sample (Ruth), ND cosmics, Reconstruction
FD Cosmics generation¶
Initial 2e6 spill sample to fulfill needs of Calibration group.
More than 4k FD cosmics declared to SAM.
Subsequent CRY files will be ran through "lite"-reco to produce PCHitLists. Processing will be done off-site via OSG resources.
FD/ND genie generation¶
An initial small sample will be produced of FD rhc/fhc nonswap/swap files and ND fhc/rhc. This will be passed to Simulation group for validation - no longer than 2 weeks. Preferred goal is to be ready for reconstruction as soon as possible.
**Week of March 3rd: Reconstruction tag. Pending changes to BadChannel interface from Jon P. Dominick to test new interface.
Waiting on completion of new gsimple files before can proceed with the genie generation, as well as picking up new physics list.
Keep-up of data through DAQ2RawDigit continues. Hope to start reconstruction when we have next tag. Database issues need to be alleviated. Stress tests of database on-going but currently hindered by two separate issues: Issue 5469 and Issue 5468
Stress tests now successful. No longer a working problem.
Adam to hand off raw2root processing to Siva et. al imminently.
Dominick to start PCHitList processing of "calibration" triggers with pending tag.
Reconstruction tag needs¶
Updated scheme for applying data-driven bad channel rates.
Finalise reconstruction modules. Including Dan's new, updated MichelEFilter? Remove Cana.
Resolve outstanding database connectivity issues for IFBeam and Validity-context based constants.
Require updated MC calibration constants before proceeding with MC reco - which relies on the MC generation step.
Production for July 2013 Argonne Collaboration Meeting¶
- Use EXISTING S12-11-16 Monte Carlo simulation input
S13-06-05 - Reconstruction
S13-06-13 - PID
S13-06-18 - CAF
- Reconstruction/PID/CAF production to be complete ideally 2 weeks prior to collaboration meeting:
Week of June 24th
- Rock and Cosmic overlays: input files are being reprocessed. - TOP PRIORITY
- Keep rock event re-use at 1:50k spills.
Complete by June 26th
*Half ND to be processed from reconstruction stage
*Time permitting full-size ND also. Full-size ND will be the guinea pig for processing using SAM.
Kanika will assist, processing MRCC
*Partial and Full FD
*Not going to be done before the July Collaboration Meeting
Action Required: Decide on partial detector size
Reconstruction Validation: Are we happy with validation now?
Reconstruction: Complete for FD beam and FD cosmics. Re-running ND overlays and starting ND full-size beam-only
PID/LEM/CAF: Complete for FD cosmics